# A physics-inspired momentum-based gradient method

### v1

@@ Abstract
In this work, a nonlinear momentum method is introduced to improve the convergence performance of momentum-based gradient optimization algorithms. The method is motivated by the dynamics of non-Newtonian mechanical systems, where conventional momentum schemes can be interpreted as a dynamical model with quadratic kinetic energy and linear damping. Based on this analogy, a generalized optimization dynamics is constructed by extending the kinetic energy formulation and incorporating a nonlinear damping term. An anharmonic kinetic energy function can be employed to represent the inertial effect of accumulated gradient information during the iterations, while the nonlinear damping mechanism enables a more flexible control of the momentum contribution along the convergence trajectory. Numerical experiments indicate that the method exhibits faster convergence and higher robustness compared to classical momentum algorithms. Moreover, its strong performance on nonconvex objectives makes it particularly suitable for inverse photonic design problems. The results suggest that dynamical systems from physics can provide a view towards the development of efficient optimization methods.
Keywords momentum method $\cdot$ optimization dynamics $\cdot$ non-Newtonian flow $\cdot$ nonlinear damping $\cdot$ inverse design
@@ 1 Introduction
Physics has long served as a fundamental source of inspiration for numerical optimization. Classical examples include Newton-type and quasi-Newton methods [1], originally derived from physical models of force balance and dynamical relaxation. Momentum-based optimization methods, such as Polyak’s Heavy Ball (HB) [2], can be interpreted as damped mechanical systems, while modern algorithms—like simulated annealing [3], Langevin dynamics [4], and approximate message passing [5]—are formulated from statistical physics principles. These physics-inspired algorithms have not only enriched optimization theory but also profoundly influenced computational physics, machine learning, and materials and chemical design.
@@
Gradient descent, one of the simplest first-order methods, has drawn much attention with the rise of machine learning, especially deep learning [6]. Its low computational cost makes it suitable for large-scale and distributed optimization. Momentum-based methods improve gradient descent by introducing a linear damping term. Its accelerated variants, including Nesterov’s Accelerated Gradient (NAG) [7] and adaptive methods such as Adam, further extend the momentum idea to improve convergence. From a physics perspective, a natural question arises: can nonlinear damping mechanisms achieve comparable or superior acceleration? Recent studies have incorporated dry friction [8], mixed viscous-Coulomb damping [9], and Hessian-dependent mechanisms, demonstrating promising performance in nonsmooth optimization.
Inverse photonicl design has emerged as a central challenge in computational physics and engineering [10]. Applications [10, 11]—from nanophotonics [12, 13] and near-field radiative heat transfer (NFRHT) to metamaterials—seek to automatically infer optimal design parameters achieving target performance. These problems are typically high-dimensional, strongly nonconvex, and computationally expensive, posing challenges for conventional optimization schemes. This motivates the exploration of physics-inspired nonlinear momentum dynamics as effective solvers for inverse design.
In this work, we develop a nonlinear momentum-based optimization algorithm and apply it to inverse design. The method accelerates convergence, enhances global search capability, and exhibits robustness under physics-constrained landscapes.
Contributions:
1. Algorithmic: We propose a gradient-based optimization method with a nonlinear momentum mechanism, inspired by non-Newtonian dynamics, which achieves faster convergence than classical momentum methods while maintaining stability.
2. Application: The algorithm is applied to inverse design in nanophotonics. Theoretical analysis and numerical experiments demonstrate its efficiency and robustness in high-dimensional, constrained problems, suggesting broader applicability to other fields such as machine learning.
@@
The remainder of this paper is organized as follows. Section 2 introduces the dynamical modeling of momentum methods and presents the nonlinear damping construction and discretization. Section 3 provides a theoretical analysis of local convergence and computational aspects. Section 4 gives numerical experiments on representative nonconvex objectives and near field thermal radiation design to validate the proposed method. Finally, Section 5 concludes the paper and outlines future works.
@@ 2.2 Anharmonic Kinetic Energy and Nonlinear Damping
The aforementioned methods are fundamentally based on linear damping and Newtonian mechanics. However, in nonconvex optimization, traditional momentum methods often exhibit large oscillations and poor stability, especially near extrema, which can result in overshooting. In many complex physical systems—such as non-Newtonian fluids, viscoelastic materials, or oscillators with nonlinear damping—the inertial and damping behaviors are not well described by linear assumptions. Recently, several studies [8, 9] have modeled optimization dynamics as nonlinear dynamical systems to tackle more challenging objectives. We expect that by introducing more sophisticated inertial or damping structures, optimization methods can achieve improved performance, particularly for nonconvex problems.
@@
Consider a finite-dimensional mechanical system with generalized coordinate $x\in\mathbb{R}^{n}$ and velocity $v=\dot{x}$. Let $K(v)$ be the kinetic energy function. Assume $K(v):\mathbb{R}^{n}\to\mathbb{R}$ is $\mathcal{C}^{2}$and strictly convex in $v$. And $V(x)$ be a $\mathcal{C}^{1}$ potential energy function; the dissipation (generalized damping) is given by a force $D(v)$ uch that $v\cdot D(v)\geq 0$for all $v$.
@@
We would like to show that the total energy decreases under these assumptions. Define the momentum:
$p:=\nabla_{v}K(v)$ (8)
The Legendre transform of $K$ is
$K^{*}(p):=\sup_{v}(p\cdot v-K(v)).$ (9)
By convex duality, the supremum is attained at the unique $v^{*}=v(p)$ satisfying $p=\nabla_{v}K(v)$, and we have the identity
$K^{*}(p)+K(v)=p\cdot v,\text{ with }v=\nabla_{p}K^{*}(p)$ (10)
Define the Hamiltonian
@@
which equals the usual energy $E=H$. The dynamics (1) can be written in Hamiltonian form with dissipation:
Thus the energy function we track is
$E(t)=H(x(t),p(t)).$ (12)
The corresponding continuous-time optimization dynamics are then described by Hamilton’s equations:
@@
If $v\cdot D(v)>0$ for every nonzero $v$, then $E$ is strictly decreasing along any trajectory with $v\neq 0$; if $v\cdot D(v)=0$ only at $v=0$, then E is nonincreasing and can decrease strictly until the system settles with $v=0$. That completes the proof under the stated assumptions. Let the dissipation force be
$D(v)=\mu|v|^{\alpha-1}\odot\text{sgn}(v)$ (19)
Where $|v|^{\alpha-1}:=(|v_{1}|^{\alpha-1},(|v_{2}|^{\alpha-1},...,|v_{n}|^{\alpha-1})^{\prime}$, and $\odot$ denotes componentwise product. Then
@@
Hence, the kenetic energy becomes,
@@
$p_{i}=\frac{\partial K}{\partial v_{i}}=\frac{1}{sm^{s-1}}s|v_{i}|^{s-1}\text{sgn}(v_{i})=m^{1-s}|v_{i}|^{s-1}\text{sgn}(v_{i})$ (22)
@@
$p_{i}=m^{1-s}|v_{i}|^{s-1}\Longrightarrow|v_{i}|^{s-1}=m^{s-1}|p_{i}|\Longrightarrow|v_{i}|=m|p_{i}|^{\frac{1}{s-1}}.$ (23)
@@
$K(v)=\frac{1}{sm^{s-1}}\sum^{n}_{i=1}|v_{i}|^{s}=\frac{1}{sm^{s-1}}\sum_{i}m^{s}|p_{i}|^{\frac{s}{s-1}}=\frac{m}{s}\sum_{i}|p_{i}|^{\frac{s}{s-1}}$ (26)
@@
$\frac{\partial K^{*}(p)}{\partial p_{i}}=\frac{m(s-1)}{s}r|p_{i}|^{r-1}\text{sgn}(p_{i})=m|p_{i}|^{r-1}\text{sgn}(p_{i}).$ (29)
But $r-1=\frac{s}{s-1}-1=\frac{1}{s-1}$, hence
$\partial_{p_{i}}K^{*}(p)=m\text{ sgn}(p_{i})|p_{i}|^{\frac{1}{s-1}}=v_{i},$ (30)
@@
$H(x,p)=K^{*}(p)+V(x)=\frac{m(s-1)}{s}\sum_{i}|p_{i}|^{r}+V(x).$ (31)
And the monotone decay of the Hamiltonian follows since $v\cdot D(v)=\geq 0$.
@@ 2.3 Discretization and Update Scheme
To apply the continuous-time dynamics to iterative optimization, we use an explicit discretization scheme. Let $x_{k}\approx x(t_{k})$ and $p_{k}\approx p(t_{k})$, the Heavy Ball method naturally arises as:
@@
here $\phi(p_{k})$ controls the damping, and $K^{*}(p_{k})$ determines the anharmonic kinetic energy. In the update scheme above, $\gamma>0$ is the nonlinear damping coefficient, $s$ controls the nonlinearity of the damping, and $h>0$ is the step size (time interval or learning rate).
The discrete updates emulate the dynamics of a velocity-dependent nonlinear dissipative system, with the following properties: when $s=2$, the algorithm reduces to the classical Heavy Ball method; when $s=1$, damping scales with the velocity magnitude, effectively suppressing high-speed oscillations; when $1<s<2$, momentum is retained in low-speed regions, helping the system escape saddle points or flat plateaus.
@@
$z_{k+1}$ $=$ $\psi(z_{k})+h\nabla K^{*}(p_{k})$ (39)
$x_{k+1}$ $=$ $\psi^{*}(x_{k})$ (40)
where $\psi(\cdot)$ is the mirror potential function and $\psi^{*}(\cdot)$ its conjuate. For example, in the $l_{p}$ geometry, $\psi(z)=\frac{1}{p}\sum|z_{i}|^{p}$ and $\psi^{*}(x)=\frac{1}{q}\sum|x_{i}|^{q}$ ( $\frac{1}{p}+\frac{1}{q}=1$).
@@ 3.1 Rosenbrock Function
The Rosenbrock function features a highly curved valley, which can be interpreted as a special potential energy landscape, making it suitable for evaluating algorithmic convergence under nonlinear coupling. In this task, we compare the following optimization algorithms: the classical Heavy Ball (HB) method, representing a momentum mechanism with linear damping and quadratic kinetic energy; the proposed nonlinear momentum method; and Nesterov’s Accelerated Gradient Decent method. As shown in Figure 1, the proposed algorithm achieves faster convergence.
Figure 1: Convergence plot of Rosenbrock function.

### v2

# A physics-inspired nonlinear momentum method for gradient descent with applications to inverse photonic design

@@ Abstract
In this work, a nonlinear momentum method is introduced to enhance the convergence performance of momentum-based gradient optimization algorithms. Classical momentum methods, such as the Heavy Ball method, can be viewed as a dynamical system with quadratic kinetic energy and linear damping. By extending this analogy to non-Newtonian dynamical systems, we construct a Hamiltonian framework for optimization problems. In this framework, nonlinear kinetic energy and nonlinear damping effects naturally emerge. It provides a more flexible and physically interpretable mechanism for optimization algorithms. Specifically, we employ an anharmonic kinetic energy function to capture the inertial effects of accumulated gradient information during the optimization process, while the nonlinear damping mechanism effectively regulates the contribution of momentum during convergence. Numerical experiments show that the proposed method achieves faster convergence compared to classical momentum algorithms, making it particularly suitable for inverse design tasks. Moreover, the Hamiltonian based algorithmic framework may offer physical insights for the development of efficient physics-inspired optimization algorithms.
Keywords momentum method $\cdot$ optimization dynamics $\cdot$ thermal photonics $\cdot$ inverse design
@@ 1 Introduction
Physics has long served as a fundamental source of inspiration for numerical optimization. Classical examples include Newton-type and quasi-Newton methods [1], which were originally derived from physical models of force balance and dynamical relaxation. Momentum-based optimization methods, such as Polyak’s Heavy Ball (HB) method [2], can be interpreted as damped dynamical systems. Modern algorithms—like simulated annealing [3], Langevin dynamics [4], and approximate message passing [5]—are formulated based on principles from statistical physics. These physics-inspired algorithms have not only enriched optimization theory and algorithms but also profoundly influenced computational physics, machine learning, and material and chemical design.
@@
Gradient descent, one of the simplest first-order methods, has drawn much attention with the rise of machine learning, especially deep learning [6]. Its low computational cost makes it suitable for large-scale and distributed optimization. Momentum based methods improve gradient descent by introducing a linear damping term [2]. Its accelerated variants, including Nesterov’s accelerated gradient descent (AGD) [7] and adaptive methods such as Adam [8], further extend the momentum idea to improve convergence. From a physics perspective, a natural question arises: can nonlinear damping mechanisms achieve comparable or superior acceleration? Recent studies have incorporated dry friction [9], mixed viscous-Coulomb damping [10], and Hessian-dependent mechanisms, demonstrating promising performance in nonsmooth optimization.
However, to the best of our knowledge, despite the significant contributions of optimization theorists, including Attouch et al. [11, 9, 10], there are still many gaps in the explicit discussion of nonlinear damping effects. Specifically, the impact of a nonlinear kinetic energy term on the performance of related gradient descent algorithms has not been explored. This paper primarily addresses these gaps by providing a Hamiltonian framework that incorporates these effects. We propose a gradient-based optimization method with a nonlinear momentum mechanism, inspired by non-Newtonian mechanics, which achieves faster convergence than classical momentum methods while maintaining stability.
Inverse design problems in photonics [12, 13, 14, 15] often rely on adjoint-based gradient computations to efficiently calculate gradients for optimization. The proposed nonlinear momentum optimization method can be integrated with these adjoint-based approaches to enhance convergence speed and stability. Specifically, the proposed method utilizes gradient information to adjust the optimization trajectory with the flexibility of nonlinear momentum, thereby accelerating convergence and improving stability. Rather than directly computing the gradients, the method uses them to update design parameters effectively, making it particularly suitable for inverse photonic design tasks. This combination not only improves the efficiency of the optimization process but also provides a novel mechanism for tackling high-dimensional, non-convex photonic design problems. Results from numerical experiments validate the efficiency of the proposed method in photonics, suggesting its broader applicability to other fields, including machine learning.
@@
The remainder of this paper is organized as follows. Section 2 introduces the dynamical modeling of momentum methods and presents the nonlinear damping construction and discretization. Section 3 provides a theoretical analysis of local convergence and computational aspects. Section 4 presents numerical experiments on representative nonconvex objectives and near-field thermal radiation design to validate the proposed method. Finally, Section 5 concludes the paper and outlines future work.
@@ 2.2 Anharmonic Kinetic Energy and Nonlinear Damping
The aforementioned methods are fundamentally based on linear damping and Newtonian mechanics. However, in the context of nonconvex optimization, these traditional frameworks often suffer from persistent oscillations and reduced stability, leading to significant overshooting near extrema. In many complex physical systems [17]—such as non-Newtonian fluids, viscoelastic materials, or oscillators with nonlinear damping—the inertial and damping behaviors are not well described by linear assumptions. Recently, several studies [9, 10] have modeled optimization dynamics as nonlinear dynamical systems to tackle more challenging objectives. Drawing inspiration from recent advances [18, 9, 10], we propose that incorporating generalized inertial and non-quadratic damping structures into gradient decent method can improve performance, particularly for nonconvex problems.
@@
Consider a finite-dimensional mechanical system with generalized coordinate $x\in\mathbb{R}^{n}$ and velocity $v=\dot{x}$. LetK(v)K(v) be the kinetic energy function. Assume $K(v):\mathbb{R}^{n}\to\mathbb{R}$ is $\mathcal{C}^{2}$and strictly convex in $v$. And $V(x)$ be a $\mathcal{C}^{1}$ potential energy function; the dissipation (generalized damping) is given by a force $D(v)$ uch that $v\cdot D(v)\geq 0$for all $v$.
@@
The dynamics (1) can be written in Hamiltonian form with dissipation, where the Hamiltonian is
@@
which equals the energy, $E(t)=H(x(t),p(t))$. The corresponding continuous-time optimization dynamics are then described by Hamilton’s equations:
@@
If $v\cdot D(v)>0$ for every nonzero $v$, then $E$ is strictly decreasing along any trajectory with $v\neq 0$, if $v\cdot D(v)=0$ only at $v=0$, then E is nonincreasing and can decrease until the system settles with $v=0$. That completes the proof under the stated assumptions. Let the dissipation force be
$D(v)=\gamma|v|^{\eta-1}\odot\text{sgn}(v)$ (15)
Where $|v|^{\eta-1}:=(|v_{1}|^{\eta-1},|v_{2}|^{\eta-1},...,|v_{n}|^{\eta-1})^{\prime}$, and $\odot$ denotes elementwise product. Then
@@
Here, we consider non-quadratic kinetic energy function,
@@
$p_{i}=\frac{\partial K}{\partial v_{i}}=\frac{m^{s-1}}{s}s|v_{i}|^{s-1}\text{sgn}(v_{i})=m^{s-1}|v_{i}|^{s-1}\text{sgn}(v_{i})$ (18)
@@
$|p_{i}|=m^{s-1}|v_{i}|^{s-1}\Longrightarrow|v_{i}|^{s-1}=m^{1-s}|p_{i}|\Longrightarrow|v_{i}|=\frac{1}{m}|p_{i}|^{\frac{1}{s-1}}.$ (19)
@@
$K(v)=\frac{m^{s-1}}{s}\sum^{n}_{i=1}|v_{i}|^{s}=\frac{m^{s-1}}{s}\sum_{i}\frac{1}{m^{s}}|p_{i}|^{\frac{s}{s-1}}=\frac{1}{s\cdot m}\sum_{i}|p_{i}|^{\frac{s}{s-1}}$ (22)
@@
$\frac{\partial K^{*}(p)}{\partial p_{i}}=\frac{(s-1)}{s\cdot m}r|p_{i}|^{r-1}\odot\text{sgn}(p_{i})=\frac{1}{m}|p_{i}|^{r-1}\odot\text{sgn}(p_{i}).$ (25)
Hence
$\partial_{p_{i}}K^{*}(p)=\frac{1}{m}\text{ sgn}(p_{i})|p_{i}|^{\frac{1}{s-1}}=v_{i},$ (26)
@@
[…]
and we previously computed
$\frac{dH}{dt}=-v\cdot D(v).$ (28)
And the monotone decay of the Hamiltonian follows since $v\cdot D(v)\geq 0$.
@@ 2.3 Discretization and Update Scheme
To apply the continuous-time dynamics to iterative optimization, we discretize the Hamilton equations. Let $x_{k}\approx x(t_{k})$, $p_{k}\approx p(t_{k})$ and set $m=1$. The Heavy Ball method naturally arises as:
@@
here $\phi(p_{k})$ controls the damping, and $K^{*}(p_{k})$ determines the anharmonic kinetic energy. In the update scheme above, $\gamma>0$ is the nonlinear damping coefficient, $\eta$ controls the nonlinearity of the damping, and $h>0$ is the step size (time interval or learning rate). The basic nonlinear momentum algorithm for minimizing $V(x)$ is given in Algorithm 1.
Algorithm 1
1:input: initial $x_{0}$, parameters $h,\gamma,\eta,r,n_{max}$
2:initialize: $p_{-1}=0$
3:for $k$ in $0:n_{max}$
[…]
6:return $x_{n_{max}}$
The discrete updates emulate the dynamics of a velocity-dependent nonlinear dissipative system, with the following properties: when $\eta=2$, the algorithm reduces to the classical Heavy Ball method; when $\eta=1$, damping scales with the velocity magnitude, effectively suppressing high-speed oscillations; when $1<\eta<2$, momentum is retained in low-speed regions, helping the system escape saddle points or flat plateaus. Furthermore, Nesterov’s AGD with nonlinear momentum for minimizing $V(x)$ is given in Algorithm 2.
Algorithm 2
1:input: initial $x_{0}$, parameters $h,\gamma,\eta,r,n_{max}$
2:initialize: $p_{-1}=0$
3:for $k$ in $0:n_{max}$
4:  $y_{k}\leftarrow x_{k}+p_{k-1}-\gamma h\phi(p_{k-1})$
5:  $p_{k}\leftarrow p_{k-1}-h\nabla V(y_{k})-h\nabla\phi(p_{k-1})$
6:  $x_{k+1}\leftarrow x_{k}+h\nabla K^{*}(p_{k})$
7:return $x_{n_{max}}$
@@
$z_{k+1}$ $=$ $\nabla\psi(x_{k})+h\nabla K^{*}(p_{k})$ (36)
$x_{k+1}$ $=$ $\nabla\psi^{*}(z_{k+1})$ (37)
where $\psi(\cdot)$ is the mirror potential function and $\psi^{*}(\cdot)$ its conjugate. For example, in the $l_{p}$ geometry, $\psi(z)=\frac{1}{p}\sum|z_{i}|^{p}$ and $\psi^{*}(x)=\frac{1}{q}\sum|x_{i}|^{q}$ ( $\frac{1}{p}+\frac{1}{q}=1$). Mirror descent with nonlinear momentum for minimizing $V(x)$ is given in Algorithm 3.
Algorithm 3
1:input: initial $x_{0}$, parameters $h,\gamma,\eta,r,q,n_{max}$
2:initialize: $p_{-1}=0$
3:for $k$ in $0:n_{max}$
4:  $p_{k}\leftarrow p_{k-1}-h\nabla V(y_{k})-h\nabla\phi(p_{k-1})$
5:  $z_{k+1}\leftarrow\nabla\psi(x_{k})+h\nabla K^{*}(p_{k})$
6:  $x_{k+1}\leftarrow\nabla\psi^{*}(z_{k+1})$
7:return $x_{n_{max}}$
3 Convergence analysis
In this section we perform a convergence analysis for the algorithm with quadratic kinetic energy with nonlinear damping. We assume objective function $V(x)$ to be $m$-strongly convex and $L$-smooth. $L$-smoothness, means that its gradient is Lipschitz continuous:
[…]
where L ¿ 0 is the Lipschitz constant.This further leads to the following descent lemma, which is fundamental in optimization:
[…]
We define the energy function,
[…]
Then, the successive energy difference can then be written as
[…]
By the descent lemma,
[…]
Substituting $v_{k+1}=x_{k+1}-x_{k}=v_{k}-h\nabla V(x_{k})-\gamma h|v_{k}|^{\eta-1}\odot\text{sgn}(v_{k})$. Let $v_{k}^{\eta-1}:=|v_{k}|^{\eta-1}\odot\text{sgn}(v_{k})$, we obtain
[…]
With the identity,
[…]
We obtain,
[…]
When $|v_{k}|<1$, $1<\eta\leq 2$, we have
[…]
If the parameters,
$(L+m)h$ $\leq$ $1$ (48)
$(1-\gamma h)^{2}-mh$ $\leq$ $0$ (49)
We get
$\Delta E\leq 0$ (50)
Based on asymptotic convergence analysis, [11] proves that the case $\eta=2$ separates weak damping ($\eta>2$) from strong damping ($\eta<2$). For a more mathematically rigorous treatment, readers may refer to this paper.
@@ 4.1 Rosenbrock Function
The Rosenbrock function features a highly curved valley, which can be interpreted as a special potential energy landscape, making it suitable for evaluating algorithmic convergence under nonlinear coupling. In this task, we compare the following optimization algorithms: the classical Heavy Ball (HB) method, representing a momentum mechanism with linear damping and quadratic kinetic energy; the proposed nonlinear momentum method; and Nesterov’s Accelerated Gradient Decent method. s Figure 1 shows for the Rosenbrock objective function, Algorithm 1 and Algorithm 2 demonstrate competitive performance against HB and AGD, outperforming them in this case.
Figure 1: Comparisons between HB, AGD, and Algorithm 1 and Algorithm 2 for the Rosenbrock objective function. The learning rate $h$ for HB and Algorithm 1 (with $\eta=s=1.9$) is 0.0002, while for AGD and Algorithm 2 (with $\eta=s=1.98$), the learning rate $h$ is 0.001. And $\gamma h=0.02$, $x_{0}=(-2.0,3.0)^{\prime}$ for all these algorithms.
Figure 2: Comparisons between different nonlinear orders for kinetic energy and dissipative potential used in Algorithm 1. The learning $h=0.0002$, damping coefficient $\gamma h=0.02$ for all these simulations.